OpenStudy (anonymous):

Translate to an equation: p $40 pants and s $20 shirts total $160 Graph the equation and use the graph to determine three different combinations of pants and shirts that total $160.

OpenStudy (jim766):

40p + 20s = 160

OpenStudy (anonymous):

160/40 = 4. 160/20 = 8. -40/20 = -2. y=-2*x+8. (4,0)(3,2)(2,4)

OpenStudy (jim766):

looks nice...great job